OpenStudy (anonymous):

An object falls downward with velocity changing according to the rule dv/dt = 400 - (v^2/16) with downward being positive and units being feet per second. a) what is the only physically realizable constant solution v sub infinity?

OpenStudy (anonymous):

b) show that if the object has initial velocity 0, then \[\lim_{t \rightarrow \infty} v(t) = v _{\infty} \] c) find the time of the fall to the ground if the object was initially at rest 2000 ft above the ground. d) what is the effect of an initial velocity of \[v _{0} = v _{\infty} + 100 \] on the time of the fall from 2000 feet?

OpenStudy (shubhamsrg):

" a) what is the only physically realizable constant solution v sub infinity?" what does this mean ?

OpenStudy (anonymous):

I think it's asking for the solution to \[v _{\infty} \] but I really am not too sure